At present, Pedro's age is 25% of his father age. Twenty years from now, Pedro's age will be 50% of his father's age. How old are they?
Answer by rothauserc(4718)

Let P be Pedro's age and F be his father's age, then
:
P = F/4
:
(P +20) = (F +20)/2
:
substitute for P in the second equation
:
(F/4 +20) = (F +20)/2
:
multiply both sides of = by 4
:
F +80 = 2F +40
:
F = 40
:
P = F/4 = 40/4 = 10
:
********************************************************
Pedro is 10 years old and Pedro's father is 40 years old
********************************************************
: